Convert the following to base ten:
a) E0T base 12
b)1011 base 2
c)43 base 5
Any help with this would be great! I am completely lost in these other number systems.

Answer by Alan3354(69443)

Convert the following to base ten:
a) E0T base 12
Don't know what that means.
----------
b)1011 base 2
= 8 + 2 + 1 = 11
------------
c)43 base 5
= 4*5 + 3 = 23